OpenStudy (anonymous):

Help The length of a side of a rectangle is 2 units less than a number, and the width is 4 units more than the length. Which of the following expressions represents the area of the rectangle, in square units? A. x^2+2x+4 B. x^2-2x+8 C. x^2-4 D. x^2+4

OpenStudy (anonymous):

It looks like it would be D but I'm not 100% sure.

OpenStudy (anonymous):

i think you are right

OpenStudy (anonymous):

I'm afraid it is not D. Here's how you get it: let x be the unknown number length is x-2 width is 4+ length, which will be 4 + x-2 = x+2 Area of a rectangle is length times width (l x w) substitute the given expression for those dimensions to calculate for area A = (length)(width) A= (x-2) (x+2) Simplify this equation or evaluate this to get correct answer.